Students in grade school were asked whether they believed Good Grades, Popularity, or Athletic Ability was the most important to them. The students responses are divided into three school locations: Rural, Suburban, or Urban. The resulting contingency table was produced from the frequencies of observations. Run a hypothesis test to determine if the goals of student are independent of the school location using α = 0.05.

Describe the dependencies if they exist.

School Area

Goals

Rural

Suburban

Urban

Grades

57

87

24

Popular

50

42

6

Sports

42

22

5
Answer #1

Chi-square test of independence

Solution:

Here, we have to use chi square test for independence of two categorical variables.

Null hypothesis: H0: The goals of student are independent of the school location.

Alternative hypothesis: Ha: The goals of student are not independent of the school location.

We are given level of significance = α = 0.05

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

We are given

Number of rows = r = 3

Number of columns = c = 3

Degrees of freedom = df = (r – 1)*(c – 1) = 2*2 = 4

α = 0.05

Critical value = 9.487729

(by using Chi square table or excel)

Calculation tables for test statistic are given as below:

Observed Frequencies

Column variable

Row variable

Rural

Suburban

Urban

Total

Grades

57

87

24

168

Popular

50

42

6

98

Sports

42

22

5

69

Total

149

151

35

335

Expected Frequencies

Column variable

Row variable

Rural

Suburban

Urban

Total

Grades

74.72239

75.72537313

17.55224

168

Popular

43.58806

44.17313433

10.23881

98

Sports

30.68955

31.10149254

7.208955

69

Total

149

151

35

335

Calculations

(O - E)

-17.7224

11.27463

6.447761

6.41194

-2.17313

-4.23881

11.31045

-9.10149

-2.20896

(O - E)^2/E

4.203332

1.678661

2.368565

0.943217

0.106909

1.754841

4.168397

2.663447

0.676864

Chi square = ∑[(O – E)^2/E] = 18.56423

P-value = 0.000957

(By using Chi square table or excel)

P-value < α = 0.05

So, we reject the null hypothesis

There is sufficient evidence to conclude that the goals of student are not independent of the school location.

There is insufficient evidence to conclude that the goals of student are independent of the school location.